# Edge-Cloud Collaborative Computing on Distributed Intelligence and Model Optimization: A Survey

### v4

@@ VII-B2 Testing Environments
Robust benchmarking of edge-cloud collaborative systems mandates controlled testing environments [243]. Specifically, accurately capturing performance characteristics requires careful configuration and control due to the complex interactions among edge devices, network infrastructure, and cloud servers. Achieving reliable benchmarking results necessitates meticulous configuration of hardware, software, and network parameters to minimize external influences and ensure reproducibility. The diverse hardware platforms in edge deployments, from resource-constrained devices to powerful edge servers, necessitate standardized testing environments for fair comparisons across different system architectures [236]. For example, BenchFaaS and KheOps employ heterogeneous testbeds, encompassing virtual machines, Raspberry Pis, and various cloud platforms, to evaluate performance under diverse hardware and network conditions [238, 239]. Additionally, accurate measurement of key performance indicators (KPIs), such as latency, throughput, and energy consumption, requires precise control over network conditions, including bandwidth, latency, and packet loss. Standardized software configurations, including operating systems, libraries, and runtime environments, are also essential for eliminating inconsistencies and ensuring comparability across experiments. Representative benchmarks that accurately reflect real-world scenarios, encompassing data volume, processing requirements, and communication patterns, are necessary for comprehensive performance evaluation.
Performance comparison frameworks offer structured methodologies for systematic evaluation of diverse ECCC architectures and optimization techniques, providing standardized approaches to performance assessment that enable identification of bottlenecks and derivation of best practices [240]. Comprehensive frameworks encompass multifaceted evaluation metrics including system performance (throughput, latency, scalability), model efficiency, quality of service, and security considerations [244], alongside standardized datasets and workloads representative of real-world applications [243]. Essential components include guidelines for establishing controlled testing environments for reproducible results [245] and tools for result analysis incorporating statistical and visualization techniques to identify performance bottlenecks and understand design trade-offs [20]. Evaluation methodologies span simulation for cost-effective scenario exploration, emulation for realistic yet controlled assessment using network emulation tools [128], and real-world testing for comprehensive validation, each offering distinct advantages in capturing system complexities while addressing specific evaluation requirements. Detailed performance comparison frameworks, systematic evaluation approaches, and comprehensive methodology analysis are provided in LABEL:sec:supp-frameworks and LABEL:sec:supp-methodologies of the Supplementary Material.
@@ VII-C System Design Evaluation
Comprehensive performance assessment across ECCC paradigms requires multi-dimensional analysis through architectural comparisons, optimization evaluations, and systematic trade-off analysis [246, 123, 247]. Architectural configurations reveal how cloud-centric solutions prioritize scalability versus edge-centric approaches like Armada emphasizing minimal latency, while hybrid systems including AppealNet and EdgeMatrix leverage predictive offloading and multi-agent algorithms for optimal balance. Additionally, optimization technique analysis examines load balancing effectiveness through two-timescale frameworks, collaborative edge-edge cooperation, and reinforcement learning-based adaptive management demonstrating superior performance in distributed scenarios [248, 249, 126]. Furthermore, trade-off analysis addresses design considerations balancing performance versus energy consumption, latency-accuracy optimization, and security-usability considerations [142, 127, 62]. Detailed architectural comparisons, optimization analysis, and trade-off methodologies are provided in LABEL:sec:supp-system-design-eval of the Supplementary Material.
@@ VIII-A Technical Challenges
As illustrated in Fig. 8, AI-driven ECCC systems face multifaceted technical challenges spanning model optimization under resource constraints, integration of heterogeneous hardware/software, performance bottlenecks, and addressing domain-specific concerns such as real-time processing and scalability. Addressing the interconnected nature of challenges requires comprehensive solutions that balance accuracy, efficiency, and scalability while ensuring seamless interoperability and overcoming practical deployment barriers.
@@ VIII-A1 Model Optimization Issues
Model optimization for edge-cloud deployment fundamentally challenges the accuracy-efficiency trade-off, where complex models demand substantial resources unsuitable for edge deployment while smaller models sacrifice accuracy [250]. Resource limitations including processing power, memory capacity, and energy availability, particularly for IoT devices [142, 251], compound deployment complexities alongside scaling challenges from communication overhead, resource contention, and device heterogeneity [252, 253, 15]. Effective optimization necessitates integrated approaches combining pruning, quantization, knowledge distillation [48] with resource-aware strategies like dynamic model selection, offloading [62, 39], energy-aware scheduling [254], and adaptive model partitioning [195]. LLMs further intensify optimization challenges, requiring specialized inference techniques for resource-constrained environments, with solutions like Shoggoth [255] demonstrating online knowledge distillation to enhance accuracy while managing data drift and computational offloading [256].
@@ VIII-A2 System Integration Challenges
System integration in ECCC demands sophisticated heterogeneity management across diverse hardware capabilities, operating systems, network bandwidths, and communication protocols between resource-constrained edge nodes and powerful cloud servers [257, 225]. Interoperability challenges stem from differing data formats, communication protocols, and software interfaces, necessitating complex data transformations and sophisticated scheduling algorithms that consider device-specific capabilities [258, 259, 128]. Standardized protocols and well-defined interfaces become essential for seamless data exchange, simplified development, and enhanced scalability, with containerization (Docker [257]) and WebAssembly [258] providing portability solutions alongside AI-driven task placement optimization [260, 259]. Standardization particularly benefits critical applications like autonomous driving [53] and smart healthcare [142] requiring reliable real-time data exchange, while promoting vendor interoperability and fostering open ecosystems [247, 261, 229].
@@ VIII-A3 Performance Bottlenecks
ECCC systems encounter significant performance bottlenecks from communication overhead introducing latency and bandwidth consumption that impacts real-time applications [116, 221, 262], while edge devices struggle with limited processing power, memory, and energy constraints when executing complex AI models and handling growing IoT data volumes [66, 263, 264, 29]. Resource limitations become particularly acute during test-time adaptation scenarios where models must adjust to new data distributions while maintaining strict latency and energy budgets [115]. Storage limitations exacerbated by device heterogeneity necessitate intelligent caching, data compression, distributed storage solutions like EdgeKV [63], and comprehensive data lifecycle management [48, 265, 266, 267]. Data distribution shifts across geographically distributed edge nodes create additional performance challenges, requiring adaptive load balancing and intelligent workload redistribution to maintain system efficiency. Real-time applications including autonomous driving, industrial automation, and smart healthcare demanding ultra-low latency [208, 268, 21] require integrated architectural innovations combining hierarchical edge computing with intelligent resource management [269], utilizing frameworks like EdgeMatrix [247] for multi-agent scaling, KubeEdge [195] for automated provisioning, predictive offloading [123], intelligent caching [270], and hardware acceleration technologies for optimal performance.
@@ VIII-A4 Practical Implementation Barriers
Real-world ECCC deployment reveals substantial barriers extending beyond theoretical optimization to encompass operational complexities, economic constraints, and integration challenges, where device heterogeneity necessitates standardized middleware layers and adaptive frameworks for automatic capability detection [20, 28, 25, 60, 49]. Network connectivity issues require robust offline operation modes and intelligent synchronization mechanisms [35], while deployment complexity often results in 40-60% timeline overruns due to integration challenges and vendor compatibility issues [226]. Economic viability concerns balance substantial upfront investments against long-term benefits, with edge computing reducing bandwidth costs by 30-50% but requiring 3-5 year payback periods [142, 45], while technical pitfalls include underestimated communication overhead [21, 253], data consistency issues requiring eventual consistency models [10], and resource contention demanding sophisticated isolation algorithms [13, 29]. Successful implementations emphasize phased rollout strategies, hybrid architectures, comprehensive security frameworks, and dedicated cross-functional teams [18, 19, 14].
@@ VIII-B Future Research Directions
Future research in ECCC must address critical scalability challenges while advancing revolutionary computing paradigms [57]. Short-term objectives include developing adaptive model partitioning algorithms [58] and standardized APIs for heterogeneous device integration [200], alongside lightweight containerization technologies optimized for resource-constrained environments. Furthermore, medium-term objectives should focus on neuromorphic computing integration [271, 272] for ultra-low-power edge inference and quantum-resistant security protocols [273, 274] for long-term protection. Additionally, long-term research directions encompass quantum edge computing [275, 276] for solving intractable optimization problems and fully autonomous edge-cloud ecosystems capable of self-optimization [259]. Consequently, implementation feasibility analysis indicates that standardization efforts represent the most immediately achievable objectives, while quantum computing integration and autonomous systems present longer-term challenges requiring significant advances in underlying technologies [17].
@@ VIII-B1 Advanced AI Techniques
Advanced AI technologies represent the next frontier in ECCC, offering transformative capabilities through intelligent automation [277] and revolutionary computational paradigms. As illustrated in Fig. 9, algorithm optimization techniques including NAS [15] and AutoML [278, 279] enable autonomous system optimization by automatically discovering optimal architectures tailored to edge-cloud constraints, thereby reducing manual configuration overhead across heterogeneous infrastructure. Furthermore, intelligent interaction capabilities encompass AI agents [260] that facilitate natural human-system interaction and autonomous decision-making for adaptive responses to dynamic conditions. Additionally, hardware & infrastructure innovations particularly neuromorphic computing [271], quantum edge computing [275], and 6G network integration [26, 280] promise revolutionary advances in computational efficiency and connectivity. Consequently, the synergistic integration of these technologies enables unprecedented system autonomy, where NAS optimizes architectures for edge constraints while AI agents coordinate workflows and quantum computing tackles intractable optimization problems, collectively transforming edge-cloud system capabilities.
@@
$\bullet$ Neural Architecture Search  NAS automates the design and optimization of neural network architectures, aiming to discover models that outperform manually designed ones in accuracy and efficiency for edge-cloud collaborative environments [281]. NAS algorithms explore a vast search space of potential architectures, evaluating their performance to identify architectures tailored to the specific constraints of edge-cloud systems, such as limited resources and bandwidth [144, 216].  Approaches like reinforcement learning, evolutionary algorithms, and gradient-based NAS show promise in finding optimal architectures for specific tasks and resource constraints [58], and combining NAS with model optimization techniques like compression and adaptation further improves performance and efficiency [282, 283].  Ultimately, integrated optimization addresses the challenges of deploying complex AI models in heterogeneous edge-cloud environments [284].
@@
$\bullet$ AutoML Automated machine learning (AutoML) streamlines AI model development and deployment in edge-cloud systems by automating processes like model selection, hyperparameter tuning, and deployment [285, 277]. Automated optimization becomes crucial in edge-cloud environments where diverse hardware and software coupled with resource constraints make manual optimization complex [286].  For example, techniques like FAST-DAD [287] distill complex models into smaller, deployable ones for edge devices, while frameworks like H2O AutoML simplify model selection and hyperparameter optimization [277].  Language-driven systems like AutoMMLab [278] further democratize AI development.  However, challenges remain in transparency, flexibility for complex scenarios, and consistent coverage of the entire ML workflow [279]. In large-scale environments, sampling-based AutoML and human-centered systems like OmniForce [288, 286] show promise, while automatically generating data-dependent configurations improves model performance [289].
@@
$\bullet$ Large Language Models Optimizing LLM inference for edge devices involves addressing limited resources and real-time performance, necessitating research on optimized inference techniques and model compression methods like Agile-Quant [290, 291].  Approaches like expert-wise bitwidth adaptation in EdgeMoE [290] can further enhance on-device efficiency.  Integrating LLMs for edge-cloud orchestration, including dynamic resource allocation and task offloading, promises more efficient systems, as demonstrated by VELO and EC-Drive [292, 207].  Developing specialized LLM architectures for edge-cloud environments, including low-rank adaptation (NetGPT [293]), model partitioning, and split learning/inference [256], is a critical research direction.
@@
$\bullet$ AI Agents Intelligent agents capable of autonomous decision-making and learning within edge-cloud systems enhance adaptability and efficiency [15, 294].  AI agents facilitate continuous learning on edge devices, addressing bandwidth limitations and connectivity issues through approaches like NeuroEvolutionary learning [282]. Collaborative learning improves model accuracy, particularly in heterogeneous networks [11, 10], and these agents optimize system performance through dynamic task offloading and resource allocation [253, 284].  Integration with lifelong learning frameworks like KubeEdge-Sedna further enhances automation [195].
@@
$\bullet$ Embodied AI Embodied AI, focusing on agents that interact with their environment, requires new algorithms and frameworks tailored for resource-constrained edge devices [295].  Algorithms for robot control and coordination must consider latency and bandwidth limitations [296], while efficient task offloading and resource allocation optimize performance [247]. Furthermore, exploring embodied AI applications for multi-robot collaboration and dynamic visual servoing presents significant research opportunities [297, 296]. Ultimately, developing robust and adaptive algorithms that effectively leverage distributed computational resources represents a critical research direction for advancing embodied AI in edge-cloud environments [298].
@@
$\bullet$ Neuromorphic Computing Inspired by the biological brain, neuromorphic computing promises energy-efficient AI models for edge-cloud environments [271, 299]. Neuromorphic hardware, like SNNs, offers potential energy savings compared to traditional hardware [272], while inherent fault tolerance improves robustness [300].  Integrating neuromorphic hardware into data centers presents challenges in algorithms, software, and benchmarks [301].  However, its potential benefits for edge robotics drive research towards ultra-low-power designs and 3D integration technologies [298, 302].
@@
$\bullet$ Quantum Edge Computing Quantum computing offers revolutionary potential for edge-cloud applications by tackling classically intractable problems [275]. Consequently, developing edge-tailored quantum algorithms [276] and robust quantum communication protocols [303], potentially in hybrid approaches, are key research avenues. However, significant security [273] and privacy concerns [274] necessitate quantum-resistant techniques. Investigating the associated societal impacts [17] is also crucial for responsible deployment.
@@
$\bullet$ 6G Network Integration 6G networks, with ultra-low latency and high bandwidth, promise to enhance edge-cloud systems [26].  Low latency enables real-time processing, while high bandwidth facilitates seamless data transfer, supporting resource-intensive AI tasks [304, 280].  Advanced features like network slicing and cell-free MIMO optimize performance [280], and integration with digital twins and pervasive AI unlocks new possibilities [305, 306].  Addressing challenges in network architecture, resource allocation, and security through AI-driven solutions is crucial [307, 27].

### v5

@@ VII-B2 Testing Environments
Robust benchmarking mandates controlled testing environments to accurately capture performance characteristics amidst complex interactions among edge devices, network infrastructure, and cloud servers [varghese2021survey]. Achieving reliable results requires meticulous configuration of hardware, software, and network parameters to minimize external influences and ensure reproducibility across diverse platforms ranging from resource-constrained devices to powerful servers [hao2019edge]. For instance, BenchFaaS [carpio2023benchfaas] and KheOps [rosendo2023kheops] employ heterogeneous testbeds encompassing virtual machines, Raspberry Pis, and cloud platforms to evaluate performance under varied conditions. Moreover, accurate measurement of KPIs such as latency and energy consumption demands precise control over network conditions, while standardized software configurations eliminate inconsistencies across experiments. Furthermore, performance comparison frameworks offer structured methodologies encompassing multifaceted evaluation metrics including system performance, model efficiency, and security considerations [bi2023openperf, yang2019frequency], alongside guidelines for controlled environments [kahlhofer2023benchmarking] and analysis tools incorporating statistical techniques [jansen2023specrg]. Additionally, evaluation methodologies spanning simulation, emulation [luckow2021exploring], and real-world testing each provide distinct advantages in capturing system complexities. Detailed frameworks and methodology analysis are provided in LABEL:sec:supp-frameworks and LABEL:sec:supp-methodologies of the Supplementary Material.
@@ VII-C System Design Evaluation
Performance assessment across ECCC paradigms requires multi-dimensional analysis through architectural comparisons, optimization evaluations, and systematic trade-off analysis [huang2021armada, li2021appealnet, ren2022edgematrix]. Architectural configurations reveal how cloud-centric solutions prioritize scalability versus edge-centric approaches like Armada emphasizing minimal latency, while hybrid systems including AppealNet and EdgeMatrix leverage predictive offloading and multi-agent algorithms for optimal balance. Additionally, optimization technique analysis examines load balancing effectiveness through two-timescale frameworks, collaborative edge-edge cooperation, and reinforcement learning-based adaptive management demonstrating superior performance in distributed scenarios [asghar2022survey, li2024twotimescale, fan2024collaborative]. Furthermore, trade-off analysis addresses design considerations balancing performance versus energy consumption, latency-accuracy optimization, and security-usability considerations [arroba2024sustainable, zhai2024edgecloud, rublein2024improved]. Detailed architectural comparisons, optimization analysis, and trade-off methodologies are provided in LABEL:sec:supp-system-design-eval of the Supplementary Material.
@@ VIII-A Technical Challenges
As illustrated in Fig. 11, the technical challenges in AI-driven ECCC can be systematically categorized into four interconnected areas that reflect the lifecycle of system design and deployment. We begin with Model Optimization Issues (Sec. VIII-A1), which are foundational challenges related to designing and adapting AI models for resource-constrained and distributed environments. Next, we examine System Integration Challenges (Sec. VIII-A2), focusing on the difficulties of creating cohesive systems from heterogeneous hardware and software. Subsequently, we analyze Performance Bottlenecks (Sec. VIII-A3), which arise during system operation due to communication overhead and resource limitations. Finally, we address Practical Implementation Barriers (Sec. VIII-A4) that emerge during real-world deployment, encompassing economic and operational complexities. Accordingly, this structure provides a logical progression from model-level design to system-level operation and deployment.
@@ VIII-A1 Model Optimization Issues
Model optimization for edge-cloud deployment fundamentally challenges the accuracy-efficiency trade-off, where complex models demand substantial resources unsuitable for edge deployment while smaller models sacrifice accuracy [gowda2024watt]. Resource limitations including processing power, memory capacity, and energy availability, particularly for IoT devices [arroba2024sustainable, ollivier2023sustainable], compound deployment complexities alongside scaling challenges from communication overhead, resource contention, and device heterogeneity [bhattacharjee2020deepedge, shi2020communicationefficient, tuli2023ai]. Effective optimization necessitates integrated approaches combining pruning, quantization, knowledge distillation [tocze2018taxonomy] with resource-aware strategies like dynamic model selection, offloading [rublein2024improved, wang2019edge], energy-aware scheduling [ullah2022framework], and adaptive model partitioning [zheng2023kubeedgesedna]. LLMs further intensify optimization challenges, requiring specialized inference techniques for resource-constrained environments, with solutions like Shoggoth [wang2023shoggoth] demonstrating online knowledge distillation to enhance accuracy while managing data drift and computational offloading [lin2024pushing].
@@ VIII-A2 System Integration Challenges
System integration in ECCC demands sophisticated heterogeneity management across diverse hardware capabilities, operating systems, network bandwidths, and communication protocols between resource-constrained edge nodes and powerful cloud servers [rac2021edge, richins2021ai]. Interoperability challenges stem from differing data formats, communication protocols, and software interfaces, necessitating complex data transformations and sophisticated scheduling algorithms that consider device-specific capabilities [menetrey2022webassembly, sowinski2023autonomous, luckow2021exploring]. Standardized protocols and well-defined interfaces become essential for seamless data exchange, simplified development, and enhanced scalability, with containerization (Docker [rac2021edge]) and WebAssembly [menetrey2022webassembly] providing portability solutions alongside AI-driven task placement optimization [maia2024survey, sowinski2023autonomous]. Standardization particularly benefits critical applications like autonomous driving [skarin2018missioncritical] and smart healthcare [arroba2024sustainable] requiring reliable real-time data exchange, while promoting vendor interoperability and fostering open ecosystems [ren2022edgematrix, shen2023edgematrix, sedlak2024equilibrium].
@@ VIII-A3 Performance Bottlenecks
ECCC systems encounter significant performance bottlenecks from communication overhead introducing latency and bandwidth consumption that impacts real-time applications [carpio2020engineering, ali-eldin2021hidden, alnoman2019emerging], while edge devices struggle with limited processing power, memory, and energy constraints when executing complex AI models and handling growing IoT data volumes [vuruma2024cloud, makaya2020costeffective, baucas2020using, kokkonen2023autonomy]. Resource limitations become particularly acute during test-time adaptation scenarios where models must adjust to new data distributions while maintaining strict latency and energy budgets [ni2024fedsl]. Storage limitations exacerbated by device heterogeneity necessitate intelligent caching, data compression, distributed storage solutions like EdgeKV [sonbol2020edgekv], and comprehensive data lifecycle management [tocze2018taxonomy, frascaria2021case, kolosov2023theory, joshi2024integration]. Data distribution shifts across geographically distributed edge nodes create additional performance challenges, requiring adaptive load balancing and intelligent workload redistribution to maintain system efficiency. Real-time applications including autonomous driving, industrial automation, and smart healthcare demanding ultra-low latency [hu2024laecips, xing2023taskoriented, xu2024fusion] require integrated architectural innovations combining hierarchical edge computing with intelligent resource management [zeng2021energyefficient], utilizing frameworks like EdgeMatrix [ren2022edgematrix] for multi-agent scaling, KubeEdge [zheng2023kubeedgesedna] for automated provisioning, predictive offloading [li2021appealnet], intelligent caching [wang2020survey], and hardware acceleration technologies for optimal performance.
@@ VIII-A4 Practical Implementation Barriers
Real-world ECCC deployment reveals substantial barriers extending beyond theoretical optimization to encompass operational complexities, economic constraints, and integration challenges, where device heterogeneity necessitates standardized middleware layers and adaptive frameworks for automatic capability detection [jansen2023specrg, kanduri2024edgecentric, bensalem2023scaling, wang2024socialized, hamm2019edge]. Network connectivity issues require robust offline operation modes and intelligent synchronization mechanisms [kawana2024communication], while deployment complexity often results in 40-60% timeline overruns due to integration challenges and vendor compatibility issues [singh2025impact]. Economic viability concerns balance substantial upfront investments against long-term benefits, with edge computing reducing bandwidth costs by 30-50% but requiring 3-5 year payback periods [arroba2024sustainable, wang2023edge], while technical pitfalls include underestimated communication overhead [xu2024fusion, shi2020communicationefficient], data consistency issues requiring eventual consistency models [liu2023distributed], and resource contention demanding sophisticated isolation algorithms [lu2024a2cdrl, kokkonen2023autonomy]. Successful implementations emphasize phased rollout strategies, hybrid architectures, comprehensive security frameworks, and dedicated cross-functional teams [poularakis2019joint, ren2021synergy, seifelnasr2024privacypreserving].
@@ VIII-B Future Research Directions
Future research in ECCC must address critical scalability challenges while advancing revolutionary computing paradigms [mungoli2023scalable]. Short-term objectives include developing adaptive model partitioning algorithms [banitalebi-dehkordi2021autosplit] and standardized APIs for heterogeneous device integration [al-aqrabi2018scalable], alongside lightweight containerization technologies optimized for resource-constrained environments. Furthermore, medium-term objectives should focus on neuromorphic computing integration [thakur2018largescale, vogginger2024neuromorphic] for ultra-low-power edge inference and quantum-resistant security protocols [baseri2024cybersecurity, lu2024quantum] for long-term protection. Additionally, long-term research directions encompass quantum edge computing [hossain2024quantumedge, nguyen2023iquantum] for solving intractable optimization problems and fully autonomous edge-cloud ecosystems capable of self-optimization [sowinski2023autonomous]. Consequently, implementation feasibility analysis indicates that standardization efforts represent the most immediately achievable objectives, while quantum computing integration and autonomous systems present longer-term challenges requiring significant advances in underlying technologies [paul2025quantumenhanced].
@@ VIII-B1 Advanced AI Techniques
Advanced AI technologies are poised to redefine the capabilities of ECCC systems. To structure this forward-looking discussion, we categorize these emerging techniques into three distinct yet complementary areas, as illustrated in Fig. 12. First, algorithm optimization focuses on automating and enhancing the AI models themselves through methods like neural architecture search and AutoML. Second, intelligent interaction explores technologies such as AI agents and embodied AI, which enable more autonomous and adaptive system behaviors. Third, hardware and infrastructure innovations, including neuromorphic computing, quantum edge computing, and 6G integration, provide the foundational advancements in compute and connectivity required to support next-generation AI. Accordingly, the following discussion details each of these promising research directions.
@@
$\bullet$ Neural Architecture Search  NAS automates the design and optimization of neural network architectures, aiming to discover models that outperform manually designed ones in accuracy and efficiency for edge-cloud collaborative environments [hao2021enabling]. NAS algorithms explore a vast search space of potential architectures, evaluating their performance to identify architectures tailored to the specific constraints of edge-cloud systems, such as limited resources and bandwidth [jain2022latencymemory, zeng2021coedge].  Approaches like reinforcement learning, evolutionary algorithms, and gradient-based NAS show promise in finding optimal architectures for specific tasks and resource constraints [banitalebi-dehkordi2021autosplit], and combining NAS with model optimization techniques like compression and adaptation further improves performance and efficiency [mannan2020clan, mih2024ecavg].  Ultimately, integrated optimization addresses the challenges of deploying complex AI models in heterogeneous edge-cloud environments [mao2024green].
@@
$\bullet$ AutoML Automated machine learning (AutoML) streamlines AI model development and deployment in edge-cloud systems by automating processes like model selection, hyperparameter tuning, and deployment [elshawi2019automated, schmitt2023automated]. Automated optimization becomes crucial in edge-cloud environments where diverse hardware and software coupled with resource constraints make manual optimization complex [xue2023omniforce].  For example, techniques like FAST-DAD [fakoor2020fast] distill complex models into smaller, deployable ones for edge devices, while frameworks like H2O AutoML simplify model selection and hyperparameter optimization [schmitt2023automated].  Language-driven systems like AutoMMLab [yang2024autommlab] further democratize AI development.  However, challenges remain in transparency, flexibility for complex scenarios, and consistent coverage of the entire ML workflow [azevedo2024multivocal]. In large-scale environments, sampling-based AutoML and human-centered systems like OmniForce [yin2024automl, xue2023omniforce] show promise, while automatically generating data-dependent configurations improves model performance [kayali2022mining].
@@
$\bullet$ Large Language Models Optimizing LLM inference for edge devices involves addressing limited resources and real-time performance, necessitating research on optimized inference techniques and model compression methods like Agile-Quant [yi2025edgemoe, shen2024agilequant].  Approaches like expert-wise bitwidth adaptation in EdgeMoE [yi2025edgemoe] can further enhance on-device efficiency.  Integrating LLMs for edge-cloud orchestration, including dynamic resource allocation and task offloading, promises more efficient systems, as demonstrated by VELO and EC-Drive [yao2024velo, chen2024edgecloud].  Developing specialized LLM architectures for edge-cloud environments, including low-rank adaptation (NetGPT [chen2024netgpt]), model partitioning, and split learning/inference [lin2024pushing], is a critical research direction.
@@
$\bullet$ AI Agents Intelligent agents capable of autonomous decision-making and learning within edge-cloud systems enhance adaptability and efficiency [tuli2023ai, wang2020convergence].  AI agents facilitate continuous learning on edge devices, addressing bandwidth limitations and connectivity issues through approaches like NeuroEvolutionary learning [mannan2020clan]. Collaborative learning improves model accuracy, particularly in heterogeneous networks [yao2023edgecloud, liu2023distributed], and these agents optimize system performance through dynamic task offloading and resource allocation [shi2020communicationefficient, mao2024green].  Integration with lifelong learning frameworks like KubeEdge-Sedna further enhances automation [zheng2023kubeedgesedna].
@@
$\bullet$ Embodied AI Embodied AI, focusing on agents that interact with their environment, requires new algorithms and frameworks tailored for resource-constrained edge devices [mu2023embodiedgpt].  Algorithms for robot control and coordination must consider latency and bandwidth limitations [tian2019fog], while efficient task offloading and resource allocation optimize performance [ren2022edgematrix]. Furthermore, exploring embodied AI applications for multi-robot collaboration and dynamic visual servoing presents significant research opportunities [tahir2023consensusbased, tian2019fog]. Ultimately, developing robust and adaptive algorithms that effectively leverage distributed computational resources represents a critical research direction for advancing embodied AI in edge-cloud environments [wan2022circuit].
@@
$\bullet$ Neuromorphic Computing Inspired by the biological brain, neuromorphic computing promises energy-efficient AI models for edge-cloud environments [thakur2018largescale, date2022neuromorphic]. Neuromorphic hardware, like SNNs, offers potential energy savings compared to traditional hardware [vogginger2024neuromorphic], while inherent fault tolerance improves robustness [rastogi2021selfrepair].  Integrating neuromorphic hardware into data centers presents challenges in algorithms, software, and benchmarks [dang2022neucasl].  However, its potential benefits for edge robotics drive research towards ultra-low-power designs and 3D integration technologies [wan2022circuit, kurshan2020case].
@@
$\bullet$ Quantum Edge Computing Quantum computing offers revolutionary potential for edge-cloud applications by tackling classically intractable problems [hossain2024quantumedge]. Consequently, developing edge-tailored quantum algorithms [nguyen2023iquantum] and robust quantum communication protocols [nguyen2024quantum], potentially in hybrid approaches, are key research avenues. However, significant security [baseri2024cybersecurity] and privacy concerns [lu2024quantum] necessitate quantum-resistant techniques. Investigating the associated societal impacts [paul2025quantumenhanced] is also crucial for responsible deployment.
@@
$\bullet$ 6G Network Integration 6G networks, with ultra-low latency and high bandwidth, promise to enhance edge-cloud systems [letaief2022edge].  Low latency enables real-time processing, while high bandwidth facilitates seamless data transfer, supporting resource-intensive AI tasks [han2021timeenergyconstrained, you20236g].  Advanced features like network slicing and cell-free MIMO optimize performance [you20236g], and integration with digital twins and pervasive AI unlocks new possibilities [tariq2022experiencedriven, baccour2023zero].  Addressing challenges in network architecture, resource allocation, and security through AI-driven solutions is crucial [adem2021how, li2021slicingbased].